OpenStudy (anonymous):

Help Please! Medal! For what values of m does the graph of y = mx^2 – 5x – 2 have no x-intercepts?

OpenStudy (tkhunny):

"m" does not mean "slope" in a quadratic equation. It is just a parameter. 1) m>0 or it will ALWAYS have intercepts. 2) x = 5/(2m) is the x-location of the minimum value. 3) Can you evaluate the function at x = 5/(2m) and observe the result.

OpenStudy (anonymous):

Here are the options: \[m>-\frac{ 25 }{ 8 }\] \[m<-\frac{ 25 }{ 8 }\] \[m<\frac{ 25 }{ 8 }\] \[m>\frac{ 25 }{ 8 }\]

OpenStudy (tkhunny):

Okay, follow the steps I provided.

OpenStudy (anonymous):

Sorry for the mistake there @Catseyeglint911

OpenStudy (anonymous):

So since m needs to be greater than zero, the best answer would be the last one. correct? ...and no worries. it's fine. ^^ @Cosmichaotic

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Okay, easier way. Can you solve: mx^2 - 5x - 2 = 0 using the Quadratic Formula?

OpenStudy (tkhunny):

BTW, my #1 was also wrong. We can have negative values. Hearty "good work" to @Cosmichaotic for making me think about it more.

OpenStudy (anonymous):

I don't think so because you don't know the value of m, and ah okay.

OpenStudy (tkhunny):

Not so! That is the key to the problem! \(b^{2} - 4ac \rightarrow (-5)^{2} - 4(m)(-2) = 25 + 8m\) I hope you recognize the discriminant from the Quadratic Formula. The deal is this, if \(25 + 8m \ge 0\), then the ARE x-intercepts. For what values of \(m\;is\;25 + 8m \lt 0\)? THAT is the question.

OpenStudy (anonymous):

-25/8 < 0

OpenStudy (anonymous):

Because m has to be a negative value that is greater than 25/8 in order to make that side of the inequality less than zero.

OpenStudy (anonymous):

Right?